Solve the system of equations.
–0.5x + 3y = –2
–0.75x – 0.5y = –0.5

Answer by checkley77(12844)

–0.5x + 3y = –2
–0.75x – 0.5y = –0.5 multiply by 6 & add
-4.5x-3y=-3
-.5x+3y=-2
---------------------
-5x=-5
x=-5/-5
x=1 answer.
-.5*1+3y=-2
-.5+3y=-2
3y=-2+.5
3y=-1.5
y=-1.5/3
y=-.5 answer.
proof:
-.75*1-.5*-.5=-.5
-.75+.25=-.5
-.5=-.5